OpenStudy (anonymous):

I keep getting the wrong number - help The standard form of the function in exercise 6 is: f(x)=3x^2+18x+15 Complete the square to write the equation in the standard form f (x) = a(x − h)^2 + k .

OpenStudy (anonymous):

I got 3(x+3)^=66

OpenStudy (henryblah):

How did you get that? Where did the f(x) go?

OpenStudy (anonymous):

f(x)=3(x+3)^2=66

OpenStudy (henryblah):

There shouldn't be an equals in the middle of your equation.